TCS new pattern Question 2013

Posted on:
tags: , , , , , , , ,

My name is PREET.But my son accidentally types the by interchanging a pair of letters in my name.What is the probability that despite this interchange, the name remains unchanged? a.5% b.20% c.25% d.12.5%
Ans:
10%

there are 10 possible pairs

out of which 1 will not change my name.

so probability that name remains unchanged = 1/10 = 10%


Roy is now 4 year older than Erik and half of that amount than Lewis.If in two years Roy will be twice as old as Erick,then in two year what would be Roy age multiplied by Lewis age??

Ans:ans is 48
firstly two equation from 1st sentence ..
suppose roy=x, erik=y, lewis=z
so x=y+4....(1)
2nd is x=z+2...(2) (Roy is now 4 year older than Erik and half of that amount than Lewis so 4/2=2)
also after 2 years erik and roy relation wil b
x+2=2(y+2).. (3)
so put this into 1st equ. u will get y=2..
and then x=6,z=4(by substituting other)
nw asking multiplication after two years of roy n lewis is (6+2)*(4+2)=48..

a man cut small cubes of 3 cubic cms each. Which he joined to form a cube wit 10 cubes length, 3 cubes in depth and 3 cubes wide. How many more small cubes does she require to form a perfect cube?

 Ans:
At present, the number of cubes used are 10*3*3=90. For a perfect cube, the cube should be 10 cubes length, 10 cubes in depth, and 10 cubes wide, i.e. 10*10*10=1000. So, additional number of cubes required is 1000-90= 910

one person has no siblings and says,"the guy in the photo is the only son of my fathers's son". what is the relation of the guy to the person

Ans:
His own son.


in a shopping mall with staff of 5 members, the average is 45.After 5 years a person joined them and the average is again 45. what is the age of 6th person?

Ans:
Since the average age is 45 yrs for the 5 members.
sum total of their ages = 45 X 5=225 yrs.
after 5 years, sum total of their ages = 225 + ( 5 X 5 ) = 250 yrs
now , let x be the age of the 6th person, so
the sum total of age is now = (250 + x ) yrs.
given the average is 45 yrs.
so,
( 250 + x )/ 6 = 45
or, x = 20 yrs.
so, the age of the 6th person is 20 years.

49 members attend the party.in that 22 are males,17 females.the shake hands between males,females,male,female.total 12 people given shake hands.how many such kinds of such shake hands are possible?

If you mean only 12 persons shook hands .
then 12C2 = 66 possible shake hands

what is the reminder when 6^17+17^6 is divided by 7?

First break the sum into two parts..
(i) 6^17 (ii) 17^6

sol-(i)- By remainder theorem there are always two remainders for any number, i.e. one is in negative and another is positive. In this case when we divide 6 by 7 we will get two remainders that is 6 and (-1). We can take any of these but for simplicity we will take (-1) is the remainder. Now (-1)^17 = -1.

sol-(ii)-- we have 17^6.. Now. 17^6 /7 = 3^6/7= (3^3)^2/7 = 27^2/7 = (-1)^2/7 = 1 is the remainder.

On adding case -(i) and (ii)... we will get final remainder = 0.

or

1   6^17 mod 7=(7-1)^17 mod 7
so 7^17 divide by 7 and remainder will 0,so we only take (-1)^17 mod 7=-1

2) 17^6 mod 7=(7*2+3)^6
so now (7*2)^17 divide by 7 and remainder will 0,so we only take (3)^6 mod 7=
729 mod 7=1
now add (1)+(2)
-1+1=0(ans)

What is the greatest possible positive integer n if 8^n divides (44)^44 without leaving a remainder?
a)14 b)28 c)29 d)15

Ans:
44^44 = 2^88 *11^44
= 8^29 * 2 * 11^44
so ans is 29

Complete the series 2, 7, 24, 77,__
Ans: 240

Ans:
(1st no.*3)+1=(2*3)+1=7
(2nd no.*3)+3=(7*3)+3=24
(3rd no.*3)+3=(24*3)+5=77
(4th no.*3)+1=(77*3)+7=238
or
ist*12=3rd 2*12=24
2nd*11=4th 7*11=77
3rd*10=5th 24*10=240
or
(2*3)+1=7
(7*3)+3=24
(24*3)+5=77
so ans is (77*3)+7=240

 

There is a circular pond and there is a road around the pond. The road is 4ft wide. The area of the pond is 11/25 of the area of the road. What is the radius of the pond?


Ans:let radius of pond=r
then,radius of road=r+4
since area of pond=11/25(area of road)
pi r^2=11/25 pi[(r+4)^2-r^2]
after solving this equa we get
25r^2-88r-176=0
r=4.94

There are two container. The volume of the second container is twice than the first. Half of the first cont. is filled with Wine & The one quarter of the second container is filled with wine. Now the remaining of the both container is filled with water. There is a third container which is enough to have the total liquid of the first two containers. What will be the ratio of the total volume and wine of the third container? Answer will be 1:3

Ans:consider A(first container)=100 for A wine filled half 100/2= 50
consider B(second container)=200 for B wine filled 1/4(200)=50
this two container mixed in 3rd container let as C=300
the ratio of wine to the volume of C is 100/300=1/3(1:3)

---------------------------------------------------------------------------------
Posted By Sundeep aka SunTechie

Sundeep is a Founder of Youth Talent Auzzar, a passionate blogger, a programmer, a developer, CISE and these days he is pursuing his graduation in Engineering with Computer Science dept.
Add Sundeep as a Friend on 


TCS New Pattern for 2012-2013 batch : mock test paper 2

Posted on:
tags: , , , ,


1. If 3y + x > 2 and x + 2y≤3, What can be said about the value of y?
A. y = -1
B. y >-1
C. y <-1
D. y = 1
Answer: B

Multiply the second equation with -1 then it will become - x - 2y≥ - 3. Add the equations. You will get y > -1.

2. If the price of an item is decreased by 10% and then increased by 10%, the net effect on the price of the item is
A. A decrease of 99%
B. No change
C. A decrease of 1%
D. An increase of 1%
Answer: C

If a certain number is increased by x% then decreased by x% or vice versa, the net change is always decrease. This change is given by a simple formula −(x10)2= −(1010)2= −1%. Negitive sign indicates decrease.

3. If m is an odd integer and n an even integer, which of the following is definitely odd?
A. (2m+n)(m-n)
B. (m+n2)+(m−n2)
C. m2+mn+n2
D. m +n
Answer: C and D (Original Answer given as D)

You just remember the following odd ± odd = even; even ± even = even; even ±odd = odd
Also odd x odd = odd; even x even = even; even x odd = even.

4. What is the sum of all even integers between 99 and 301?
A. 40000
B. 20000
C. 40400
D. 20200
Answer: D

The first even number after 99 is 100 and last even number below 301 is 300. We have to find the sum of even numbers from 100 to 300. i.e., 100 + 102 + 104 + ............... 300.
Take 2 Common. 2 x ( 50 + 51 + ...........150)
There are total 101 terms in this series. So formula for the sum of n terms when first term and last term is known is n2(a+l)
So 50 + 51 + ...........150 = 1012(50+150)
So 2 x 1012(50+150) = 20200

5. There are 20 balls which are red, blue or green. If 7 balls are green and the sum of red balls and green balls is less than 13, at most how many red balls are there?
A. 4
B. 5
C. 6
D. 7
Answer: B

Given R + B + G = 17; G = 7; and R + G < 13. Substituting G = 7 in the last equation, We get R < 6. So maximum value of R = 6

6. If n is the sum of two consecutive odd integers and less than 100, what is greatest possibility of n?
A. 98
B. 94
C. 96
D. 99
Answer : C

We take two odd numbers as (2n + 1) and (2n - 1).
Their sum should be less than 100. So (2n + 1) + (2n - 1) < 100 ⇒ 4n < 100.
The largest 4 multiple which is less than 100 is 96

7. x2 < 1/100, and x < 0 what is the highest range in which x can lie?
A. -1/10 < x < 0
B. -1 < x < 0
C. -1/10 < x < 1/10
D. -1/10 < x
Answer: A

Remember:
(x - a)(x - b) < 0 then value of x lies in between a and b.
(x - a)(x - b) > 0 then value of x does not lie inbetween a and b. or ( −∞, a) and (b, −∞) if a < b
x2 < 1/100 ⇒
(x2−1/100)<0⇒(x2−(1/10)2)<0⇒(x−1/10)(x+1/10)<0
So x should lie inbetween - 1/10 and 1/10. But it was given that x is -ve. So x lies in -1/10 to 0

8. There are 4 boxes colored red, yellow, green and blue. If 2 boxes are selected, how many combinations are there for at least one green box or one red box to be selected?
A. 1
B . 6
C. 9
D. 5
Answer: 5

Total ways of selecting two boxes out of 4 is 4C2 = 6. Now, the number of ways of selecting two boxes where none of the green or red box included is only 1 way. (we select yellow and blue in only one way). If we substract this number from total ways we get 5 ways.

9. All faces of a cube with an eight - meter edge are painted red. If the cube is cut into smaller cubes with a two - meter edge, how many of the two meter cubes have paint on exactly one face?
A. 24
B. 36
C. 60
D. 48
Answer : A

If there are n cubes lie on an edge, then total number of cubes with one side painting is given by 6×(n−2)2. Here side of the bigger cube is 8, and small cube is 2. So there are 4 cubes lie on an edge. Hence answer = 24

10. Two cyclists begin training on an oval racecourse at the same time. The professional cyclist completes each lap in 4 minutes; the novice takes 6 minutes to complete each lap. How many minutes after the start will both cyclists pass at exactly the same spot where they began to cycle?
A. 10
B. 8
C. 14
D. 12
Answer: D

The faster cyclyst comes to the starting point for every 4 min so his times are 4, 8, 12, ......... The slower cyclist comes to the starting point for every 6 min so his times are 6, 12, 18, ......... So both comes at the end of the 12th min.

11. M, N, O and P are all different individuals; M is the daughter of N; N is the son of O; O is the father of P; Among the following statements, which one is true?
A. M is the daughter of P
B. If B is the daughter of N, then M and B are sisters
C. If C is the granddaughter of O, then C and M are sisters
D. P and N are bothers.
Answer: B

From the diagram it is clear that If B is the daughter of N, then M and B are sisters. Rectangle indicates Male, and Oval indicates Female.

12. In the adjoining diagram, ABCD and EFGH are squres of side 1 unit such that they intersect in a square of diagonal length (CE) = 1/2. The total area covered by the squares is

A. Cannot be found from the information
B. 1 1/2
C. 1 7/8
D. None of these
Answer: C


Let CG = x then using pythogerous theoremCG2+GE2=CE2
⇒ x2+x2=(1/2)2⇒2x2=1/4⇒x2=1/8
Total area covered by two bigger squares = ABCD + EFGE - Area of small square = 2 - 1/8 = 15/8

13. There are 10 stepping stones numbered 1 to 10 as shown at the side. A fly jumps from the first stone as follows; Every minute it jumps to the 4th stone from where it started - that is from 1st it would go to 5th and from 5th it would go to 9th and from 9th it would go to 3rd etc. Where would the fly be at the 60th minute if it starts at 1?
A. 1
B. 5
C. 4
D. 9
Answer : A

Assume these steps are in circular fashion.
Then the fly jumps are denoted in the diagram. It is clear that fly came to the 1st position after 5th minute. So again it will be at 1st position after 10th 15th .....60th. min.

So the fly will be at 1st stone after 60th min.



14. What is the remainder when 617+1176 is divided by 7?
A. 1
B. 6
C. 0
D. 3
Answer: C

617 = (7−1)17 =
17C0.717−17C1.716.11.....+17C16.71.116−17C17.117

If we divide this expansion except the last term each term gives a remainder 0. Last term gives a remainder of - 1.

Now From Fermat little theorem, [ap−1p]Rem=1
So [1767]Rem=1

Adding these two remainders we get the final remainder = 0

15. In base 7, a number is written only using the digits 0, 1, 2, .....6. The number 135 in base 7 is 1 x 72 + 3 x 7 + 5 = 75 in base 10. What is the sum of the base 7 numbers 1234 and 6543 in base 7.
A. 11101
B. 11110
C. 10111
D. 11011
Answer: B


In base 7 there is no 7. So to write 7 we use 10. for 8 we use 11...... for 13 we use 16, for 14 we use 20 and so on.
So from the column d, 4 + 3 = 7 = 10, we write 0 and 1 carried over. now 1 + 3 + 4 = 8 = 11, then we write 1 and 1 carried over. again 1 + 2 + 5 = 8 = 11 and so on

16. The sequence {An} is defined by A1 = 2 and An+1=An+2n what is the value of A100
A. 9902
B. 9900
C. 10100
D. 9904
Answer: A
We know that A1 = 2 so A2=A1+1=A1+2(1)=4
A3=A2+1=A2+2(2)=8
A4=A3+1=A3+2(3)=14
So the first few terms are 2, 4, 8, 14, 22, ......
The differences of the above terms are 2, 4, 6, 8, 10...
and the differences of differences are 2, 2, 2, 2. all are equal. so this series represents a quadratic equation.
Assume An = an2+bn+c
Now A1 = a + b + c = 2
A2 = 4a + 2b + c = 4
A3 = 9a + 3b + c = 8
Solving above equations we get a = 1, b = - 1 and C = 2
So substituting in An = n2+bn+c = n2−n+2
Substitute 100 in the above equation we get 9902.

17.Find the number of rectangles from the adjoining figure (A square is also considered a rectangle)

A. 864
B. 3276
C. 1638
D. None

Answer: C

To form a rectangle we need two horizontal lines and two vertical lines. Here there are 13 vertical lines and 7 horizontal lines. The number of ways of selecting 2 lines from 13 vertical lines is 13C2 and the number of ways of selecting 2 lines from 7 horizontals is 7C2. So total rectangles = 7C2x13C2

18. A, B, C and D go for a picnic. When A stands on a weighing machine, B also climbs on, and the weight shown was 132 kg. When B stands, C also climbs on, and the machine shows 130 kg. Similarly the weight of C and D is found as 102 kg and that of B and D is 116 kg. What is D's weight
A. 58kg
B. 78 kg
C. 44 kg
D. None
Answer : C

Given A + B = 132; B + C = 130; C + D = 102, B + D = 116
Eliminate B from 2nd and 4th equation and solving this equation and 3rd we get D value as 44.

19. Roy is now 4 years older than Erik and half of that amount older than Iris. If in 2 years, roy will be twice as old as Erik, then in 2 years what would be Roy's age multiplied by Iris's age?
A. 28
B. 48
C. 50
D. 52
Answer: 48

20. X, Y, X and W are integers. The expression X - Y - Z is even and the expression Y - Z - W is odd. If X is even what must be true?
A. W must be odd
B. Y - Z must be odd
C. W must be odd
D. Z must be odd
Answer: A or C (But go for C)

21. Mr and Mrs Smith have invited 9 of their friends and their spouses for a party at the Waikiki Beach resort. They stand for a group photograph. If Mr Smith never stands next to Mrs Smith (as he says they are always together otherwise). How many ways the group can be arranged in a row for the photograph?
A. 20!
B. 19! + 18!
C. 18 x 19!
D. 2 x 19!
Answer: C

22. In a rectanglular coordinate system, what is the area of a triangle whose vertices whose vertices have the coordinates (4,0), (6, 3) adn (6 , -3)
A. 6
B. 7
C. 7.5
D. 6.5
Answer: A

23. A drawer holds 4 red hats and 4 blue hats. What is the probability of getting exactly three red hats or exactly three blue hats when taking out 4 hats randomly out of the drawer and immediately returning every hat to the drawer before taking out the next?
A. 1/2
B. 1/8
C. 1/4
D. 3/8
Answer: B

24. In how many ways can we distribute 10 identical looking pencils to 4 students so that each student gets at least one pencil?
A. 5040
B. 210
C. 84
D. None of these
Answer: C

25. The prime factorization of intezer N is A x A x B x C, where A, B and C are all distinct prine intezers. How many factors does N have?
A. 12
B. 24
C. 4
D. 6
Answer: A

26. Tim and Elan are 90 km from each other.they start to move each other simultanously tim at speed 10 and elan 5 kmph. If every hour they double their speed what is the distance that Tim will pass until he meet Elan
A. 45
B. 60
C. 20
D. 80
Answer: B

27. A father purchases dress for his three daughter. The dresses are of same color but of different size .the dress is kept in dark room .What is the probability that all the three will not choose their own dress.
A. 2/3
B. 1/3
C. 1/6
D. 1/9
Answer: B

28. N is an integer and N>2, at most how many integers among N + 2, N + 3, N + 4, N + 5, N + 6, and N + 7 are prime integers?
A. 1
B. 3
C. 2
D. 4
Answer: C

29. A turtle is crossing a field. What is the total distance (in meters) passed by turtle? Consider the following two statements
(X) The average speed of the turtle is 2 meters per minute
(Y) Had the turtle walked 1 meter per minute faster than his average speed it would have finished 40 minutes earlier
A. Statement X alone is enough to get the answer
B. Both statements X and Y are needed to get the answer
C. Statement Y alone is enough to get the answer
D. Data inadequate
Answer: B

30. Given the following information, who is youngest?
C is younger than A; A is talled than B
C is older than B; C is younger than D
B is taller than C; A is older than D
A. D
B. B
C. C
D. A
Answer: B

31. If P(x) = ax4+bx3+cx2+dx+e has roots at x = 1, 2, 3, 4 and P(0) = 48, what is P(5)
A. 48
B. 24
C. 0
D. 50
Answer: A

DISCLAIMER: Please note that this Mock test is intended to be supplementary and cannot substitute thorough preparation. Questions in this test are simulated based on TCS Aptitude Test pattern as of August 2012. However, test patterns and questions are likely to change anytime without notice.

All The Best !  


---------------------------------------------------------------------------------
Posted By Sundeep aka SunTechie

Sundeep is a Founder of Youth Talent Auzzar, a passionate blogger, a programmer, a developer, CISE and these days he is pursuing his graduation in Engineering with Computer Science dept.
Add Sundeep as a Friend on 

TCS New Questions Pattern for 2012-13 batch: Moct Test Paper 1

Posted on:
tags: , , , ,
TCS MOCK TEST 1

TCS New Questions Pattern for 2012-13 batch: Moct Test Paper 1

TCS changed their aptitude test patterns but still questions comes from old pattern...
So, all 70 questions pattern I posted on my blog so that my readers should prepare for both old as well as new pattern...

Click here for old pattern questions 


1.The difference between the ages of two of my three grandchildren is 3. My eldest grandchild is three times older than the age of my youngest grandchild and my eldest grandchild's age is two years more than the ages of my two youngest grandchildren added together. How old is my eldest grandchild?
(a) 12
(b) 13
(c) 10
(d) 15

2.A greengrocer was selling apple at a penny each, chickoos at 2 for a penny and peanuts at 3 for a penny. A father spent 7 pennies and got the same amount of each type of fruit for each of his three children. What did each child get?
(a) 1 apple, 2 chickoos, 2 peanuts (c) 1 apple, 3 chickoos, 2 peanuts
(b) 1 apple, 2 chickoos, 1 peanut (d) 1 apple, 1 chickoo, 1 peanut

3.The IT giant Tirnop has recently crossed a head count of 150000 and earnings of $7 billion. As one of the forerunners in the technology front, Tirnop continues to lead the way in products and services in India. At Tirnop, all programmers are equal in every respect. They receive identical salaries and also write code at the same rate.Suppose 12 such programmers take 12 minutes to write 12 lines of code in total. How long will it take 72 programmers to write 72 lines of code in total?
(a) 12
(b) 18
(c) 6
(d) 72

4.One day Rapunzel meets Dwarf and Byte in the Forest of forgetfulness. She knows that Dwarf lies on Mondays, Tuesdays and Wednesdays, and tells the truth on the other days of the week. Byte, on the other hand, lies on Thursdays, Fridays and Saturdays, but tells the truth on the other days of the week. Now they make the following statements to Rapunzel - Dwarf: Yesterday was one of those days when I lie. Byte: Yesterday was one of those days when I lie too. What day is it?
(a) Monday
(b) Sunday
(c) Thursday
(d) Saturday

5.Alok and Bhanu play the following min-max game. Given the expression N = 9 + X + Y – Z , where X, Y and Z are variables representing single digits (0 to 9). Alok would like to maximize N while Bhanu would like to minimize it. Towards this end, Alok chooses a single digit number and Bhanu substitutes this for a variable of her choice (X, Y or Z). Alok then chooses the next value and Bhanu, the variable to substitute the value. Finally Alok proposes the value for the remaining variable. Assuming both play to their optimal strategies, the value of N at the end of the game would be
(a) 20
(b) 18
(c) 27
(d) 0

6.A sheet of paper has statements numbered from 1 to 40. For each value of n from 1 to 40, statement n says "At least n of the statements on this sheet are true." Which statements are true and which are false?
a) All statements are false. b) The odd numbered statements are true the even numbered are false. c) All statements are true. d) The even numbered statements are true and the odd numbered are false.

7.10 suspects are rounded by the police and questioned about a bank robbery. Only one of them is guilty. The suspects are made to stand in a line and each person declares that the person next to him on his right is guilty. The rightmost person is not questioned. Which of the following possibilities are true?
(A) All suspects are lying or the leftmost suspect is innocent. (B) All suspects are lying and the leftmost suspect is innocent. (c) Both A and B (d) Neither A nor B

8.The citizens of planet nigiet are 8 fingered and have thus developed their decimal system in base 8. A certain street in nigiet contains 1000 (in base 8) buildings numbered 1 to 1000. How many 3s are used in numbering these buildings?
(a) 192
(b) 64
(c) 54
(d) 102

9.On planet zorba, a solar blast has melted the ice caps on its equator. 8 years after the ice melts, tiny plantoids called echina start growing on the rocks. echina grows in the form of a circle and the relationship between the diameter of this circle and the age of echina is given by the formula d = 4 * √ (t - 8) for t ≥ 8 where d represents the diameter in mm and t the number of years since the solar blast. Jagan recorded the radius of some echina at a particular spot as 8mm. How many years back did the solar blast occur?
(a) 8
(b) 12
(c) 16
(d) 24

10.A circular dartboard of radius 1 foot is at a distance of 20 feet from you. You throw a dart at it and it hits the dartboard at some point Q in the circle. What is the probability that Q is closer to the center of the circle than the periphery?
(a) ¼
(b) ½
(c)3/4
(d) 1/3

11.After the typist writes 12 letters and addresses 12 envelopes, she inserts the letters randomly into the envelopes (1 letter per envelope). What is the probability that exactly 1 letter is inserted in an improper envelope?
(a) 11/12
(b) 0
(c) 1/12
(d) 1/6

12.Alok is attending a workshop "How to do more with less" and today's theme is “Working with fewer digits”. The speakers discuss how a lot of miraculous mathematics can be achieved if mankind (as well as womankind) had only worked with fewer digits. The problem posed at the end of the workshop is “How many 5-digit numbers can be formed using the digits 1, 2, 3, 4, 5 (but with repetition) that are divisible by 4?” Can you help Alok find the answer?
(a) 375
(b) 625
(c) 500
(d) 3125

13.Given 3 lines in the plane such that the points of intersection form a triangle with sides of length 20, 20 and 30, the number of points equidistant from all the 3 lines is
(a) 1
(b) 0
(c) 4
(d) 2

14.The pacelength P is the distance between the rear of two consecutive footprints. For men, the formula, n/P = 144 gives an approximate relationship between n and P where, n = number of steps per minute and P = pacelength in meters. Bernard knows his pacelength is 164cm. The formula applies to Bernard’s walking. Calculate Bernard’s walking speed in kmph. (a) 236.16 (b) 11.39 (c) 8.78 (d) 23.24

15.Alice and Bob play the following coins-on-a-stack game. 20 coins are stacked one above the other. One of them is a special (gold) coin and the rest are ordinary coins. The goal is to bring the gold coin to the top by repeatedly moving the topmost coin to another position in the stack. Alice starts and the players take turns. A turn consists of moving the coin on the top to a position i below the top coin (0 ≤ i ≤ 20). We will call this an i-move (thus a 0-move implies doing nothing). The proviso is that an i-move cannot be repeated; for example once a player makes a 2-move, on subsequent turns neither player can make a 2-move. If the gold coin happens to be on top when it's a player's turn then the player wins the game. Initially, the gold coin is the third coin from the top. Then
(a) In order to win, Alice’s first move should be a 0-move. (b) In order to win, Alice’s first move should be a 1-move. (c) Alice has no winning strategy. (d) In order to win, Alice’s first move can be a 0-move or a 1-move

16.For the FIFA world cup, Paul, the octopus has been predicting the winner of each match with amazing success. It is rumored that in a match between 2 teams A and B, Paul picks A with the same probability as A's chances of winning. Let's assume such rumors to be true and that in a match between Ghana and Bolivia, Ghana the stronger team has a probability of 2/3 of winning the game. What is the probability that Paul will correctly pick the winner of the Ghana-Bolivia game?
(a) 5/9
(b) 1/9
(c) 2/3
(d) 1/3

17.36 people {a1, a2, ..., a36} meet and shake hands in a circular fashion. In other words, there are totally 36 handshakes involving the pairs, {a1, a2}, {a2, a3}, ..., {a35, a36}, {a36, a1}. Then size of the smallest set of people such that the rest have shaken hands with at least one person in the set is
(a) 18
(b) 13
(c) 34
(d) 12

18.One the Planet Oz, there are 8 days in a week – Sunday to Saturday and another day called Oz day. There are 36 hours in a day and each hour has 90 minutes while each minute has 60 seconds. As on earth, the hour hand covers the dial twice every day. Find the approximate angle between the hands of clock on Oz when the time is 12.40 am.
(a) 89
(b) 251
(c) 111
(d) 79

19.The IT giant Tirnop has recently crossed a head count of 150000 and earning of $7 billion. As one of the forerunners in the technology front, Tirnop continues to lead the way in products and services in India. At Tirnop, all programmers are equal in every respect. They receive identical salaries and also write code at the same rate. If 12 such programmers take 12 minutes to write 12 lines of code in total, how many lines of code can be written by 72 programmers in 72 minutes?
(a) 6
(b) 432
(c) 72
(d) 12

20.A hollow cube of size 5 cm is taken with a thickness of 1 cm. It is made of smaller cubes of size 1 cm. If the 4 faces of the outer surface of the cube are painted totally, how many faces of the smaller cubes remain unpainted?
(a) 900
(b) 488
(c) 500
(d)800

21.Planet fourfi resides in 4-dimensional space and thus the currency used by its residents are 3-dimensional objects. The notes are cubical in shape while their coins are spherical. However the coin minting machinery lays out some stipulations on the size of the coins. - The diameter of the coins should be at least 64mm and not exceed 512mm. - Given a coin, the diameter of the next larger coin is at least 50% greater. - The diameter of the coin must always be an integer. You are asked to design a set of coins of different diameters with these requirements and your goal is to design as many coins as possible. How many coins can you design?
(a) 5
(b) 8
(c) 9
(d) 6

22.A hare and a tortoise race along a circle of 100 yards diameter. The tortoise goes in one direction and the hare in the other. The hare starts after tortoise has covered 1/5 of its distance and that too leisurely. The hare and tortoise meet when the hare has covered only 1/8 of the distance. By what factor should the hare increase its speed so as to tie the race?
(a)8
(b) 37.80
(c) 40
(d) 5

23.Anoop managed to draw 7 circles of equal radii with their centres on the diagonal of a square such that the two extreme circles touch two sides of the square and each middle circle touches two circles on either side. Find the ratio of the radius of the circles to the side of the square.
(a) 1:(2 + 7Ö2)
(b) 1:(4 + 7Ö3)
(c) (2 + 7Ö2):1
(d) 1:(2 + 6Ö2)

24.Ferrari S.p.A. is an Italian sports car manufacturer based in Maranello, Italy. Founded by Enzo Ferrari in 1928 as Scuderia Ferrari, the company sponsored drivers and manufactured race cars before moving into production of street-legal vehicles in1947 as Ferrari S.p.A. Throughout its history the company has been noted for its continued participation in racing especially in Formula One where it has enjoyed great success. Rohit once bought a Ferrari. It could go 2 times as fast as Mohit’s old Mercedes. If the speed of Mohit’s Mercedes is 32 km/hr and the distance travelled by the Ferrari is 952 km, find the total time taken in hours for Rohit to drive that distance.
(a) 15.88
(b) 29.75
(c) 14.88
(d)476

25.There are two boxes, one containing 10 red balls and the other containing 10 green balls. You are allowed to move the balls between the boxes so that when you choose a box at random and a ball at random from the chosen box, the probability of getting a red is maximized. This maximum probability is
(a) 37/38
(b) 1 / 2
(c) 14/19
(d) 3 / 4

26.Pizza shops make pizzas of same thickness but different diameter. Cost of pizza A with diameter 8 cm is 80 $, cost of the pizza B with diameter 12 cm is 240 $, cost of the pizza B with diameter 24 cm is 720 $. Which of the above mentioned pizzas gives the best value for money?
(a) A
(b) B
(c) C
(d) Cannot say


27.Determine the distance between x-intercept and z-intercept of the plane whose equation is 6x+8y-3z=72.
(a) 31.92
(b)26.83
(c) 32.66
(d) 25.63

28.Lucy finds around 25 groups of stars that appear to her as constellations. She draws 7 patterns of the constellations in her notebook and notes down the number of stars in each of them. She counts 5 stars in first constellation and 15 on next. She counts a number the third time and forgets to note it down. The next four constellations she counts 51, 53, 159, 161. Next day her father looks at the notebook and wants to know the number of stars in the third constellation. Lucy only remembers that number of stars counted in each of the constellation followed a pattern 5, 15, x, 51, 53, 159, 161. What is the value of x?
(a) 19
(b) 17
(c) 47
(d) 31

29.6 persons standing in the queue for ROBERT movie are wearing different coloured shirts. All of them belong to different age groups. After two years, their average age will be 43. A seventh person joined with them, hence the current average age has become 45. Find the age of seventh person?
(a) 67
(b) 69
(c) 72
(d) 74

30.X is 6 years younger to Y. X's father is a businessman who invested 10000 at 8% rate of interest and obtained his amount after 10 years. Y's father is a job holder who invested around 20000 at 2% rate and obtained his amount after 20 years. Now compounding, both of them get around Rs. A. After 5 years, the ratio of ages of X and Y is 1:2. Now X's father is 20 years older to Y and Y's father is 30 years more than X. After 20 years, again X's mother asks X's father to purchase a LCD TV which costs around 45000. What is the age of X and Y together?
(a) 12
(b) 8
(c) 18
(d) 6

31.The great musician Rahman has organized a live concert. The concert is organized in a big auditorium. Rahman plays both English and Tamil songs on his Yamaha Casio. The audience in the eastern part of the auditorium love listening to Tamil songs and those in the western part of the auditorium love listening to English songs. He plays songs in random. The probability that he plays English songs for 6 consecutive times is 1 in
(a) 32
(b) 16
(c) 64
(d) 128

32.It is dark in my bedroom and I want to get two socks of the same color from my drawer, which contains 24 red and 24 blue socks. How many socks do I have to take from the drawer to get at least two socks of the same colour?
(a) 3
(b) 25
(c) 48
(d) 26

33.A person was fined for exceeding the speed limit by 10mph. Another person was also fined for exceeding the same speed limit by twice the same. If the second person was traveling at a speed of 35 mph, find the speed limit
(a) 35 mph
(b) 15 mph
(c) 20 mph
(d) 30 mph

34.A person drives with constant speed and after some time he sees a milestone with 2 digits. Then travels for 1 hour and sees the same 2 digits in reverse order. 1 hour later he sees that the milestone has the same 2 digits with a 0 between them. What is the speed of the car?
(a) 54 (b) 45 (c) 27 (d) 36 35. With four fifths of the tank full, a vehicle travels 12 miles. How much distance will the vehicle travel with one-third tank full? (a) 8.05 km (b) 6.05 km (c) 12 km (d) 5 km
DISCLAIMER: Please note that this Mock test is intended to be supplementary and cannot substitute thorough preparation. Questions in this test are simulated based on TCS Aptitude Test pattern as of August 2012. However, test patterns and questions are likely to change anytime without notice.

All The Best !  

---------------------------------------------------------------------------------
Posted By Sundeep aka SunTechie

Sundeep is a Founder of Youth Talent Auzzar, a passionate blogger, a programmer, a developer, CISE and these days he is pursuing his graduation in Engineering with Computer Science dept.
Add Sundeep as a Friend on 
◘ >